# Trust Me, I’m Wrong: High-Certainty Hallucinations in LLMs

### v1

@@ 3.4 Models and Datasets
We evaluate CHOKE prevalence on TriviaQA (Joshi et al., 2017) and Natural Questions (Kwiatkowski et al., 2019), two common English closed-book question-answering datasets. We use three base models and their instruction-tuned versions: Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Llama-3.1-8B, Llama-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B, Gemma-2-27B and Gemma-2-9B-it (Team et al., 2024). See Appendix A for details regarding the evaluation of each model and setting-specific dataset.
@@ 4 Results
First, we identify whether CHOKE examples are consistently present across certainty metrics and across different models (Section 4.1). Next, we show that CHOKE examples are consistently present in instruct-tuned and larger models (Section 4.2). Lastly, we show that CHOKE can not be explained as noise (Section 4.3).
4.1 CHOKE Consistently Exists
As shown in Figure 3, a non-negligible amount of hallucinations occur with high certainty, demonstrating the existence of CHOKE. For example, for the Mistral model with the probability certainty measure on the child setting (upper-left subfigure of Figure 3), approximately $20\%$ of hallucinations exceed the certainty threshold ($T^{*}=0.5$). This means that using $0.5$ as the certainty threshold, a significant portion of hallucinations would be classified as certain hallucinations. Similar patterns are found in other setups shown in the figure.
These findings highlight that high-certainty hallucinations are not rare but a common phenomenon in these models. See Appendix B for a qualitative evaluation showing examples of hallucinations that exhibit high probability and low semantic entropy scores, both measures of high certainty.
CHOKE is Consistent Across Certainty Thresholds.
The black dashed line in each subfigure represents the optimal certainty threshold (explained in Section 3.3). Our results show that between $10\%$ and $40\%$ of hallucinations exceed this threshold, confirming the presence of certain hallucinations. The figure shows that this trend persists across a range of certainty thresholds, not just the optimal one, demonstrating that the results are robust to threshold selection.
While the correct answers (blue line) are consistently above the hallucinations (red line) across all thresholds, the certainty-correct relationship is not absolute. Many correct answer samples occur with low certainty, indicating that certainty levels vary even for correct predictions while the model knows the answer. See Appendix D for similar results on the Gemma and Llama models. Additionally, results using a temperature of $0.5$ instead of $1$ for semantic entropy generations are provided in Appendix E. For similar results using a variation of the child setting, demonstrating the robustness of the phenomenon, refer to Appendix F.
CHOKE is Consistent Across Certainty Metrics.
As the figure shows, the extent of high-certainty hallucinations varies across the certainty measures (different rows). Nevertheless, across all three measures, we effectively identify CHOKE examples.
4.2 CHOKE Persists in Instruction-Tuning and Larger Models
Since instruction tuning and model size often influence model behavior, we investigate their effect on CHOKE to assess its persistence.
@@ 4.2 CHOKE Persists in Instruction-Tuning and Larger Models
Instruction-tuned models display poorer calibration between uncertainty and hallucinations, as reflected by their higher rates of high-certainty hallucinations. For example, for the Mistral model with the probability certainty measure on the child setting, approximately $20\%$ of hallucinations have a certainty value of $\mathbf{0.5}$ or higher. In contrast, for the Mistral-Instruct model, around $20\%$ of hallucinations have a certainty value of $\mathbf{0.9}$ or higher. These results are shown in the upper-left subfigure and the upper-third-from-the-left subfigure of Figure 3. Similar results are found with Llama and Gemma models (Appendix D).
@@
(a) Gemma, Child
(b) Gemma, Alice-Bob
Figure 4: Detection of CHOKE: Comparing Gemma-27B to Gemma-9B on TriviaQA. The results indicate that certainty levels are comparable and slightly higher for the larger model, Gemma-27B.
CHOKE also appears in larger models.
@@
Next, we conduct the same test on the larger Gemma-2-27B using TriviaQA in both settings. The results are shown in Figure 4. Evidently, the certainty levels of the Gemma-2-27B hallucinations are comparable to, and in some cases slightly higher than, those observed in Gemma-9B. This suggests that not only does this phenomenon exist in larger models, but it may also become more pronounced as model size increases. See Appendix H for similar results on the Natural Questions dataset.
Probability Semantic Entropy Probability Difference
Model Dataset Random Certain Random Certain Random Certain
Llama TriviaQA 3.55 27.42 2.67 16.67 2.21 18.29
NQ 4.38 21.21 3.03 7.14 4.57 16.02
Mistral TriviaQA 5.11 28.21 3.93 15.66 4.99 22.97
NQ 12.51 34.53 6.08 15.14 12.72 31.75
Gemma TriviaQA 3.46 22.99 3.39 14.29 1.83 17.24
NQ 5.74 26.52 3.71 9.2 5.6 22.11
Llama-Inst TriviaQA 7.33 19.41 4.01 15.77 6.58 18.05
NQ 7.25 18.08 4.59 15.36 6.82 17.34
Mistral-Inst TriviaQA 8.03 36.48 7.26 19.05 7.24 34.49
NQ 11.05 31.46 11.29 24.94 8.65 30.83
Gemma- Inst TriviaQA 8.07 35.73 8.16 27.01 7.46 32.01
NQ 7.6 31.72 8.03 25.42 10.08 35.47
Table 2: Jaccard Similarity of CHOKE across different prompts. The Certain columns shows the overall similarity of CHOKE samples between prompts in the TriviaQA and NaturalQA datasets, using Probability, Semantic entropy, and Probability Difference as the certainty thresholds. Results indicate high similarity, suggesting consistency across settings. All scores are statistically significant (permutation test, Random column, $p<0.0001$ for probability and probability difference, $p<0.0007$ for semantic entropy).
4.3 CHOKE Cannot Be Explained as Noise
@@
While the existence of CHOKE is apparent, a potential criticism is that these samples could merely reflect noise stemming from the natural correlation between uncertainty and hallucinations, rather than constituting a distinct and consistent subset. To address this, we evaluate the similarity of CHOKE examples across two distinct contexts: the Alice-Bob and Child settings.
@@
Hallucination and non-hallucination classifications differ significantly between these settings, with overall Jaccard similarity between their hallucinations ranging from 30% to 50%. Thus, finding consistent CHOKE samples across these diverse settings will suggest they are not artifacts of the uncertainty-hallucination correlation but instead represent a robust phenomenon.
@@
We quantify this consistency using the Jaccard similarity of CHOKE across settings and validate its uniqueness with a permutation test on $10$K randomly sampled subsets of hallucination samples of equivalent size. The results confirm that CHOKE similarity between context settings exceeds random expectations, as can be seen in Table 2, using probabilities, semantic entropy, and probability difference certainty metrics. Appendix G provides additional analyses, including results for only shared hallucination examples between the settings that further demonstrate the uniqueness of CHOKE.
(a) TriviaQA, Child
(b) Natural Questions, Alice-Bob
Figure 5: Failure of Certainty-Based Mitigation Methods: Percentage of unmitigated hallucinations (Y-axis) across six models (X-axis) for three mitigation methods: Sampling, Predictive Entropy, and Probability. Lower values indicate higher success in mitigation. All methods fail to address a significant portion of CHOKE, likely due to their dependence on uncertainty scores.
5 Hallucination Mitigation and CHOKE
Having established the existence of CHOKE, we hypothesize that they may impact the performance of mitigation methods. To test this hypothesis, we experiment with hallucination mitigation while focusing on CHOKE.
5.1 Mitigation Based on Certainty Measures
Recent research has proposed leveraging certainty measures to mitigate hallucinations, primarily by abstaining from generating outputs when the model is uncertain about its predictions Cole et al. (2023); Tjandra et al. (2024). These approaches often rely on probabilities and entropy-based metrics Tjandra et al. (2024) or combine them with self-evaluation techniques Tomani et al. (2024). Other techniques employ alternative information-theoretic measures Yadkori et al. (2024a, b) or involve cross-verification using multiple language models to assess uncertainty Feng et al. (2024a).
The effectiveness of these measures is often evaluated using the area under the receiver operator characteristic (AUROC), which measures how well uncertainty ranks correct versus incorrect responses Kuhn et al. (2023). However, this focus on ranking accuracy overlooks cases where hallucinations occur with high certainty or when correct answers are generated with low certainty. Therefore, this oversight may limit their ability to mitigate hallucinations comprehensively.
To evaluate whether CHOKE samples are being overlooked, we test the usefulness of uncertainty-based mitigation methods in our setting. Specifically, we use the approaches of Tomani et al. (2024) and Cole et al. (2023), incorporating three uncertainty measures they used: negative log-likelihood, sampling, and predictive entropy. We introduce the probability measure as a simplified alternative to negative log-likelihood. Below, we briefly describe each measure, See Appendix C.3 for details.
Probability is a simplified version of negative log-likelihood that considers only the probability of the first token. Sampling assess the diversity of the model’s generated outputs, under the assumption that greater diversity reflects lower certainty. Lastly, predictive entropy estimates uncertainty by evaluating the average unpredictability of the model’s outputs. Examples are mitigated only if their uncertainty is below a defined threshold, as described in Section 3.3.
5.2 Certainty-based Mitigation Methods Fail on CHOKE
The results in Figure 5 show the percentage of unmitigated hallucination examples for each model. A lower value is better, as it indicates fewer hallucinations remain unmitigated.
Our results in Figure 5 show that a significant proportion of hallucinations exceed the threshold across all methods and models, remaining unmitigated. Among the methods, sampling performs the worst, while probability and predictive entropy consistently show a lower percentage of unmitigated hallucinations. This suggests sampling is less effective in avoiding mislabeling hallucinations among uncertainty-based methods. However, even probability and predictive entropy fail to mitigate a notable percentage of hallucinations. Lowering the threshold could improve mitigation, but this would lead to an increased mislabeling of low-certainty correct answers, as shown in Section 4.
These findings show that certainty-based mitigation fails to address high-certainty hallucinations. While uncertainty-based methods are expected to struggle with high-certainty hallucinations, our results here and in Section 4 indicate they consistently miss a distinct type, regardless of the uncertainty metric or threshold strictness. This failure suggests a fundamental limitation in assuming an ideal uncertainty metric can fully capture hallucinations.

### v2

@@ 3.4 Models and Datasets
We evaluate CHOKE prevalence on TriviaQA (Joshi et al., 2017) and Natural Questions (Kwiatkowski et al., 2019), two common English closed-book question-answering datasets.
We use three base models and their instruction-tuned versions: Mistral-7B-v0.3, Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Llama-3.1-8B, Llama-3.1-8B-Instruct (Dubey et al., 2024), Gemma-2-9B, Gemma-2-27B and Gemma-2-9B-it (Team et al., 2024). Unless stated otherwise, the Gemma model we use is Gemma-2-9B. Details in Appendix A.
To maintain readability, results in the tables and figures in the main body are based on Natural Questions; similar results on TriviaQA are given in the appendix and referred to in relevant sections. Figures show results on different prompt settings, while similar results on other prompts are referred to in the relevant sections.
Certainty Method Llama Mistral Gemma Llama-It Mistral-It Gemma-It
Probability $17.2_{\scriptscriptstyle\pm 2.7}$ $39.6_{\scriptscriptstyle\pm 2.4}$ $20.1_{\scriptscriptstyle\pm 2.1}$ $30.1_{\scriptscriptstyle\pm 4.3}$ $28.7_{\scriptscriptstyle\pm 2.8}$ $28.7_{\scriptscriptstyle\pm 5.3}$
Probability Diff. $17.2_{\scriptscriptstyle\pm 2.6}$ $42.1_{\scriptscriptstyle\pm 4.6}$ $19.5_{\scriptscriptstyle\pm 4.1}$ $26.7_{\scriptscriptstyle\pm 3.2}$ $27.2_{\scriptscriptstyle\pm 2.1}$ $29.0_{\scriptscriptstyle\pm 3.8}$
Semantic Entropy $17.9_{\scriptscriptstyle\pm 5.3}$ $20.0_{\scriptscriptstyle\pm 2.0}$ $15.8_{\scriptscriptstyle\pm 4.1}$ $19.7_{\scriptscriptstyle\pm 4.4}$ $31.2_{\scriptscriptstyle\pm 3.7}$ $24.0_{\scriptscriptstyle\pm 2.6}$
Metrics Intersection $5.87_{\scriptscriptstyle\pm 1.07}$ $7.96_{\scriptscriptstyle\pm 1.57}$ $6.38_{\scriptscriptstyle\pm 0.91}$ $9.08_{\scriptscriptstyle\pm 1.45}$ $15.48_{\scriptscriptstyle\pm 1.31}$ $12.55_{\scriptscriptstyle\pm 2.34}$
Table 1: Percentages of CHOKE hallucinations. CHOKE examples occur in 16–43% of hallucinations outputs across models (‘it’ short for instruct) and certainty methods. Key finding: Many hallucinations occur with high certainty, showing models can produce confident hallucination responses even when they have the correct knowledge.
MistralMistral-Instruct
Figure 3: Analysis of CHOKE across thresholds. Cumulative distributions of hallucinations (H) and correct answers (NH) when models possess correct knowledge. The X-axis shows the certainty; The Y-axis shows cumulative percentages. Black dashed lines mark optimal certainty thresholds for separating hallucinations from correct answers.
@@ 4 Results
First, we identify whether CHOKE examples are consistently present across certainty metrics and across different models (Section 4.1). Next, we show that CHOKE examples are consistently present in instruct-tuned and larger models (Section 4.2). Lastly, we show that CHOKE examples cannot be explained as noise (Section 4.3).
4.1 CHOKE Examples Are Widespread
[…]
See similar results on TriviaQA in Appendix D. For qualitative evaluation that shows certain hallucinations across models, see Appendix B.
CHOKE examples persist across certainty metrics.
As the table shows, the extent of high-certainty hallucinations varies across the certainty measures (different rows). Nevertheless, across all three measures, we effectively identify CHOKE examples. In addition, the intersection between the certainty measures ranges from 5% to 16%, showing that while each method may highlight different aspects of uncertainty (Beigi et al., 2024), CHOKE examples are not artifacts of a specific metric, and they persist even under strict agreement.
CHOKE examples persist across certainty thresholds.
Figure 3 shows an evaluation of certain hallucination percentages for any given threshold. The black dashed line in each subfigure represents the optimal certainty threshold (explained in Section 3.3). Our results show that between $16\%$ and $43\%$ of hallucinations exceed this threshold, confirming the presence of certain hallucinations. The figure shows that this trend persists across a range of certainty thresholds, not just the optimal one, showing that the results are robust to threshold selection.
While the correct answers (blue line) are consistently above the hallucinations (red line) across all thresholds, the certainty-correct relationship is not absolute. Many correct answer samples occur with low certainty, indicating that certainty levels vary even for correct predictions while the model knows the answer. See Appendix D for similar results on the Gemma and Llama models, on additional prompts, and on TriviaQA. Similar results with temperature of $0.5$ instead of $1$ for semantic entropy generations are in Appendix E.
Semantic Entropy Probability
Model Random CHOKE Random CHOKE
Llama $5.2{\pm 1.8}$ $\mathbf{15.8}{\pm 6.6}$ $5.1{\pm 1.4}$ $\mathbf{25.7}{\pm 12.8}$
Mistral $6.4{\pm 1.0}$ $\mathbf{18.2}{\pm 3.9}$ $13.6{\pm 2.5}$ $\mathbf{40.6}{\pm 12.1}$
Gemma $4.3{\pm 1.3}$ $\mathbf{13.0}{\pm 5.5}$ $5.7{\pm 1.5}$ $\mathbf{27.8}{\pm 15.5}$
Llama-Inst $5.3{\pm 1.4}$ $\mathbf{18.0}{\pm 7.0}$ $8.6{\pm 2.2}$ $\mathbf{25.8}{\pm 11.5}$
Mistral-Inst $10.2{\pm 2.4}$ $\mathbf{24.9}{\pm 9.2}$ $9.3{\pm 2.1}$ $\mathbf{31.2}{\pm 13.4}$
Gemma-Inst $7.7{\pm 2.4}$ $\mathbf{26.7}{\pm 12.5}$ $9.2{\pm 2.1}$ $\mathbf{31.7}{\pm 16.6}$
Table 2: CHOKE examples are more consistent across prompts than other hallucinations. The CHOKE columns show high Jaccard similarity across prompts, indicating strong consistency, while randomly sampled hallucinations (Random) have low similarity. All results are statistically significant (permutation test, $p<0.008$).
4.2 CHOKE Examples Persists in Instruction-Tuning and Larger Models
Since instruction tuning and model size often influence model behavior, we investigate their effect on CHOKE examples to assess their persistence.
@@ 4.2 CHOKE Examples Persists in Instruction-Tuning and Larger Models
Instruction-tuned models display poorer calibration between uncertainty and hallucinations, as reflected by Figure 3. For example, for the Mistral model with the probability measure, approximately $20\%$ of hallucinations have a certainty value of $\mathbf{0.5}$ or higher. In contrast, for the Mistral-Instruct model, around $20\%$ of hallucinations have a certainty value of $\mathbf{0.9}$ or higher. Similar results are found with Llama and Gemma models (Appendix D).
@@
Figure 4: Detection of CHOKE: Comparing Gemma-27B to Gemma-9B on hallucination (H) and non-hallucination (NH) data, shows similar certainty levels.
CHOKE examples also appear in larger models.
@@
Next, we conduct the same test on the larger Gemma-2-27B. The results are shown in Figure 4. Evidently, the certainty levels of the Gemma-2-27B hallucinations are comparable to those observed in Gemma-9B. This suggests that this phenomenon also exist in larger models. See Appendix H for similar results on the TriviaQA dataset.
4.3 CHOKE Examples Cannot Be Explained as Noise
@@
While the existence of CHOKE examples is apparent, a potential criticism is that these samples could merely reflect noise stemming from the natural correlation between uncertainty and hallucinations, rather than constituting a distinct and consistent subset. To address this, we evaluate the similarity of CHOKE examples across any two of our prompt variants.
@@
Hallucination and non-hallucination classifications differ significantly between these settings, with overall Jaccard similarity between their hallucinations ranging from 30% to 50%. Thus, finding consistent CHOKE examples across these diverse settings will suggest they are not artifacts of the uncertainty-hallucination correlation but instead represent a robust phenomenon.
@@
We quantify this consistency using the Jaccard similarity of CHOKE examples across settings and validate its uniqueness with a permutation test on 1010K randomly sampled subsets of hallucination samples of equivalent size. [^3] We ran this for any two settings and report the mean. The results confirm that CHOKE examples similarity between context settings exceeds random expectations, as shown in Table 2, using semantic entropy and probability metrics. Appendix G provides additional analyses, including results for TriviaQA and for only shared hallucination examples between the settings that further demonstrate the uniqueness of CHOKE.
(a) CHOKE-Score for different mitigation methods.
(b) CHOKE-Score vs. standard metrics.
Figure 5: Our mitigation outperforms other methods on CHOKE-Score and CHOKE-Score reveals limits of standard methods. Figure 5(a) (left) shows our probe method achieves the highest CHOKE-Score scores. Figure 5(b) (right) compares CHOKE-Score (red) to other metrics (blue shades), showing that certainty-based methods perform well generally but poorly on CHOKE-Score, exposing gaps in handling CHOKE examples. Probe methods maintain more consistent performance, demonstrating stronger robustness. Scores averaged over six models and all prompts.
5 CHOKE-Score
Having established the existence and distinctiveness of CHOKE examples, we hypothesize that the performance of hallucination mitigation methods will differ when evaluated on CHOKE examples, compared to their overall effectiveness. These cases represent a unique challenge, and assessing mitigation on them may expose limitations or trade-offs not captured by standard metrics. To this end, we introduce CHOKE-Score, a novel evaluation metric for assessing the ability of hallucination mitigation methods to reduce CHOKE examples. This perspective is particularly valuable in high-stakes settings, where model certainty impacts decisions.
Formally, given CHOKE examples detected via a certainty measurement method $d$ (Section 3.2), we measures the proportion successfully mitigated:
$\text{CHK-d}_{(\mathcal{M},\mathcal{C}_{d})}$ $=\frac{|\mathcal{M}\cap\mathcal{C}_{d}|}{|\mathcal{C}_{d}|}$ (2)
Here, $\mathcal{C}_{d}$ is the set of CHOKE examples flagged by a detection method $d$, and $\mathcal{M}$ is the set of successfully mitigated hallucinations. The score ranges from 0 to $1$, with $1$ indicating that all $\mathcal{C}_{d}$ examples were mitigated (found in $\mathcal{M}$). To ensure a broad definition of CHOKE-Score, we combine all three detection methods from Section 3.2 to create two score variations:
$\text{CHK}_{(\mathcal{M},\mathcal{C}_{\cap})}$ $=\frac{|\mathcal{M}\cap\mathcal{C}_{\cap}|}{|\mathcal{C}_{\cap}|}$ (3)
$\text{CHK-F}_{(\mathcal{M},\mathcal{C}_{\cup})}$ $=\frac{|\mathcal{M}\cap\mathcal{C}_{\cup}|}{|\mathcal{C}_{\cup}|}$ (4)
where $\mathcal{C}_{\cap}$ is the set of CHOKE examples flagged by all detection methods (intersection), and $\mathcal{C}_{\cup}$ is the set flagged by at least one method (union). These two variants allow us to evaluate mitigation effectiveness under strict (CHK) and flexible (CHK-F) detection criteria.
While CHK targets strict cases flagged by all methods, CHK-F includes any case flagged by at least one method. Together, they offer a fuller picture of mitigation performance.
5.1 Mitigation Methods
We use the CHOKE-Score to evaluate three hallucination mitigation approaches: certainty-based, prompt-based, and probe-based, including our own variant: CHOKE-tuned probe method. [^4] When required, training uses a 50%/50% random split.
5.1.1 Baseline Methods
Certainty methods.
Certainty-based mitigation leverages uncertainty estimates to determine when to abstain from generation. Such approaches rely on self-evaluation techniques Tomani et al. (2024), information-theoretic measures Yadkori et al. (2024a, b), or use cross-verification across multiple models to assess uncertainty Feng et al. (2024). We employ a sampling-based method following Cole et al. (2023) and a predictive entropy-based method following Tomani et al. (2024).
Prompting methods.
Prompt-based methods aim to improve factuality or guide the model toward abstention or generating truthful responses by using crafted prompts (Feng et al., 2024; Taveekitworachai et al., 2024). Specifically, we use a self-reflect prompt Feng et al. (2024) and null-shot prompting Taveekitworachai et al. (2024). As these methods require instructions, we evaluate them only on instruction-tuned models.
Probing methods.
Probe-based methods use classifiers trained on internal activations to detect and abstain from hallucinations (Li et al., 2023; CH-Wang et al., 2023). We apply logistic regression on residual stream activations at the last token of the 15th layer, following CH-Wang et al. (2023).
5.1.2 CHOKE-tuned Mitigation (Ours)
We augment the standard linear probe by oversampling CHOKE examples during training, specifically increasing their proportion in training to 65%. We hypothesize this will boost CHOKE-Score performance with minimal impact on overall accuracy.
5.2 Results
We report both CHOKE-Score variants—CHK (strict) and CHK-F (flexible)—for each mitigation. For context, we include overall accuracy (ACC), hallucination accuracy (H-ACC), and non-hallucination accuracy (NH-ACC). Figure 5 shows the results averaged across six models. [^5] Results are on NaturalQA. Similar results on TriviaQA and additional metrics are in Appendix I.
CHOKE examples challenge mitigation. Figure 5(a) compares the performance of the different mitigation methods on the CHOKE-Score. Prompting methods perform worst, followed by certainty-based mitigation. Probe-based methods perform better, and our CHOKE-Tuned performs best. This demonstrates that focusing on CHOKE examples may help improve on CHOKE-Score, an important consideration in domains where reliable, high-certainty predictions are essential. However, while the score increased, fully mitigating high-certainty hallucinations despite having the correct knowledge remains challenging.
CHOKE-Score reveals a hidden performance gap. Figure 5(b) presents a clear trend: methods strong on general metrics often underperform on CHOKE. While certainty-based methods achieve high accuracy overall and on general hallucinations, their CHOKE-Score is much lower. Prompt-based methods perform well on some metrics yet show the largest drop in CHOKE-Score. In contrast, linear probe mitigation remains stable across all metrics, with only a minor gap on CHOKE-Score.
These findings support our hypothesis that CHOKE examples differ from general hallucinations and require separate evaluation. Traditional metrics often overlook this, hiding methods’ weaknesses on CHOKE examples. CHOKE-Score fills this gap, offering a more accurate picture of a method’s robustness, especially where model certainty matters. Our CHOKE-tuned mitigation boosts performance, showing that targeted training can enhance robustness on CHOKE examples.